OpenStudy (anonymous):

Given that x-3 is a factor of the expression x^3-5x^2+kx-15, show that the value of k is 11.......need help.....

OpenStudy (anonymous):

have you heard of the remainder theorem?

OpenStudy (anonymous):

well yes......but i couldn't understand the question....

OpenStudy (anonymous):

what you do is to substitute 3 (not -3 because you need the 'opposite') into the equation. \[(3)^3-5(3)^2+k(3)-15\] now since x-3 is a factor this expresssion ^^ will equal 0 (because there is no remainder) solve to find d

OpenStudy (anonymous):

*k not d

OpenStudy (anonymous):

ok...so is it....9-45+3k-15...........and then?

OpenStudy (anonymous):

@prasheelasingh make that equal to 0 and then solve for k

OpenStudy (anonymous):

k=17?

OpenStudy (anonymous):

\[\frac{ 2x+6 }{ 5 }-\frac{x+7}{3}<\frac{3x8}{2}\] help needed please.......

OpenStudy (anonymous):

\[<\frac{3x-8}{2}\] correction....

OpenStudy (anonymous):

@prasheelasingh the first 3 is cubed, not squared